complete the ordered pairs for the equation:
2x + y = 10.
(5, ), ( , 10), ( , -2), (0, ).

Answer by checkley75(3666)

2X+Y=10
2*5+Y=10
10+Y=10
Y=10-10
Y=0 (5,0)
-------------
2X+10=10
2X=10-10
X=0 (0,10)
-------------
2X-2=10
2X=12
X=12/2
X=6 (6,-2)
--------------
2*0+Y=10
0+Y=10
Y=10 (0,10)
--------------